# Zihan Tan    Suyuan Huang    Guancheng Wan    Wenke Huang    He Li    Mang Ye

### v3

@@ Abstract
Federated Graph Learning (FGL) combines the privacy-preserving capabilities of federated learning (FL) with the strong graph modeling capability of Graph Neural Networks (GNNs). Current research addresses subgraph-FL from the structural perspective, neglecting the propagation of graph signals on spatial and spectral domains of the structure. From a spatial perspective, subgraph-FL introduces edge disconnections between clients, leading to disruptions in label signals and a degradation in the semantic knowledge of the global GNN. From a spectral perspective, spectral heterogeneity causes inconsistencies in signal frequencies across subgraphs, which makes local GNNs overfit the local signal propagation schemes. As a result, spectral client drift occurs, undermining global generalizability. To tackle the challenges, we propose a global knowledge repository to mitigate the challenge of poor semantic knowledge caused by label signal disruption. Furthermore, we design a frequency alignment to address spectral client drift. The combination of Spatial and Spectral strategies forms our framework $S^{2}$FGL. Extensive experiments on multiple datasets demonstrate the superiority of $S^{2}$FGL. The code is available at https://github.com/Wonder7racer/S2FGL.git
@@ 1 Introduction
Figure 1: In the first place, compared with centralized GNN training, subgraph-FL is encountering label signal disruption challenges, leading to decreased structure inertia score and poor semantic knowledge for GNNs. Moreover, we demonstrate the heat map of Kullback-Leibler divergence of eigenvalue distributions across clients. Inconsistency in subgraph signal frequency caused by spectral heterogeneity leads to spectral client drift.
Figure 2: Illustration of the problems. (a) From the spatial perspective, nodes in subgraph-FL lose label signals from originally nearby labeled nodes due to edge loss, namely label signal disruption. Correspondingly, GNNs suffer from poor semantic knowledge, leading to a deteriorated global GNN. (b) From the spectral perspective, spectral heterogeneity induces inconsistencies in signal frequencies across subgraphs, leading to spectral client drift in the signal propagation paradigms of GNNs and degraded global generalizability.
@@ 2 Related Work
Federated Learning. Federated learning (Huang et al., 2023c, 2024; Yang et al., 2023; Wan et al., 2024a) has gained increasing attention in recent years as it addresses the issue of data silos while ensuring data privacy. Several research directions have emerged from FL, including robustness (Xu et al., 2022; Hong et al., 2023; Zhu et al., 2023; Fang & Ye, 2022), fairness (Chen et al., 2024; Ezzeldin et al., 2023; Ray Chaudhury et al., 2022), and asynchronous federated learning (Xu et al., 2023; Zhang et al., 2023d). Generally, FL can be categorized into two main types by their optimization objective: traditional FL (tFL) and personalized FL (Hu et al., 2024; Shang et al., 2022; Lv et al., 2024; Smith et al., 2017). Research of tFL aims at aggregating a highly generalizable global model (McMahan et al., 2017; Li et al., 2020; Acar et al., 2021; Zhang et al., 2022a). For instance, FedNTD (Lee et al., 2022) preserves the global perspective on local data for the not-true classes, FEDGEN (Zhu et al., 2021) ensembles user information in a data-free manner to regulate local training, SCAFFOLD (Karimireddy et al., 2020) uses variance reduction for the client drift phenomenon. Instead, strategies of pFL aim to customize models that perform optimally for each client (Wu et al., 2023; Zhou & Konukoglu, 2023; Li et al., 2021; Zhang et al., 2023b). Specifically, FedALA (Zhang et al., 2023c) proposed adaptive masks to achieve personalized aggregation, DBE (Zhang et al., 2023a) stores domain biases for eliminating them, FedRoD (Chen & Chao, 2022) leverages two heads for global and personalized tasks.
@@ 4.1 Motivation
Motivation of NLIR. Graph data is fragmented across multiple clients in subgraph-FL, which inevitably disrupts semantic signals from labeled nodes across clients. Label signal disruption leads to key pathways for propagating semantic information, resulting in biased local feature representations that ultimately degrade GNN ability. For validation, we investigate the relationship between the decrease in SIS and the client scale. Specifically, the SIS exhibits a downward trend as the client scale increases, with notably lower scores in the range emphasized by mainstream subgraph-FL studies, thereby highlighting the label signal disruption phenomenon. We aim to mitigate its negative impact by preserving valuable semantic knowledge across fragmented subgraphs. Correspondingly, we propose Node Label Information Reinforcement (NLIR). By selecting nodes with both structural representativeness and rich semantic information for the construction of a global repository and injecting it during local training, NLIR reduces the information loss inherent in the subgraph scenarios. Subsequently, NLIR assesses the similarity distributions between node features and all representative prototypes for both local and global GNNs during local training, thereby enabling multidimensional semantic localization of nodes. By aligning the two similarity distributions, it effectively injects semantic knowledge and enhances feature modeling semantically.

### v4

@@ Abstract
Federated Graph Learning (FGL) combines the privacy-preserving capabilities of Federated Learning (FL) with the strong graph modeling capability of Graph Neural Networks (GNNs). Current research addresses subgraph-FL from the structural perspective, neglecting the propagation of graph signals on the spatial and spectral domains of the structure. From a spatial perspective, subgraph-FL introduces edge disconnections between clients, leading to disruptions in label signals and a degradation in the semantic knowledge of the global GNN. From a spectral perspective, spectral heterogeneity causes inconsistencies in signal frequencies across subgraphs, which makes local GNNs overfit the local signal propagation schemes. As a result, spectral client drift occurs, undermining global generalizability. To tackle the challenges, we propose a global knowledge repository to mitigate the challenge of poor semantic knowledge caused by label signal disruption. Furthermore, we design a frequency alignment to address spectral client drift. The combination of Spatial and Spectral strategies forms our framework $S^{2}$FGL. Extensive experiments on multiple datasets demonstrate the superiority of $S^{2}$FGL. The code is available at https://github.com/Wonder7racer/S2FGL.git
@@ 1 Introduction
Figure 1: In the first place, compared with centralized GNN training, subgraph-FL is encountering label signal disruption challenge, leading to decreased structure inertia score and poor semantic knowledge for GNNs. Moreover, we demonstrate the heat map of the Kullback-Leibler divergence of eigenvalue distributions across clients. Inconsistency in subgraph signal frequency caused by spectral heterogeneity leads to spectral client drift.
Figure 2: Problem Illustration. (a) From the spatial perspective, nodes in subgraph-FL lose label signals from originally nearby labeled nodes due to edge loss, namely label signal disruption. Correspondingly, GNNs suffer from poor semantic knowledge, leading to a deteriorated global GNN. (b) From the spectral perspective, spectral heterogeneity induces inconsistencies in signal frequencies across subgraphs, leading to spectral client drift in the signal propagation paradigms of GNNs and degraded global generalizability.
@@ 2 Related Work
Federated Learning. Federated learning (Huang et al., 2023c, 2024; Yang et al., 2023; Wan et al., 2024a) has gained increasing attention in recent years as it addresses the issue of data silos while ensuring data privacy. Several research directions have emerged from FL, including robustness (Xu et al., 2022; Hong et al., 2023; Zhu et al., 2023; Fang & Ye, 2022), fairness (Chen et al., 2024; Ezzeldin et al., 2023; Ray Chaudhury et al., 2022), and asynchronous federated learning (Xu et al., 2023; Zhang et al., 2023d). Generally, FL can be categorized into two main types by their optimization objective: traditional FL (tFL) and personalized FL (Hu et al., 2024; Shang et al., 2022; Lv et al., 2024; Smith et al., 2017). Research of tFL aims at aggregating a highly generalizable global model (McMahan et al., 2017; Li et al., 2020; Acar et al., 2021; Zhang et al., 2022a). For instance, FedNTD (Lee et al., 2022) preserves the global perspective on local data for the not-true classes, FEDGEN (Zhu et al., 2021) ensembles user information in a data-free manner to regulate local training, and SCAFFOLD (Karimireddy et al., 2020) uses variance reduction for the client drift phenomenon. Instead, strategies of personalized FL (pFL) aim to customize models that perform optimally for each client (Wu et al., 2023; Zhou & Konukoglu, 2023; Li et al., 2021; Zhang et al., 2023b). Specifically, FedALA (Zhang et al., 2023c) proposed adaptive masks to achieve personalized aggregation, DBE (Zhang et al., 2023a) stores domain biases for elimination, and FedRoD (Chen & Chao, 2022) leverages two heads for global and personalized tasks.
@@ 4.1 Motivation
Motivation of NLIR. Graph data is fragmented across clients in subgraph-FL, which inevitably disrupts semantic signals from labeled nodes across clients. Label signal disruption undermines key pathways for propagating semantic information. This results in biased local feature representations, which ultimately degrade the performance of GNNs. For validation, we investigate the relationship between the decrease in SIS and the client scale. Specifically, the SIS exhibits a downward trend as the client scale increases, with notably lower scores in the range emphasized by mainstream subgraph-FL studies, thereby highlighting the label signal disruption phenomenon. We aim to mitigate its negative impact by preserving valuable semantic knowledge across fragmented subgraphs. Correspondingly, we propose NLIR. By selecting nodes with both structural representativeness and rich semantic information for the construction of a global repository and injecting it during local training, NLIR reduces the information loss inherent in the subgraph scenarios. Subsequently, NLIR assesses the similarity distributions between node features and all representative prototypes for both local and global GNNs during local training, thereby enabling multidimensional semantic localization of nodes. By aligning the two similarity distributions, it effectively injects semantic knowledge and enhances feature modeling semantically.